# A Neural Material Point Method for Particle-based Simulations

### v1

@@ Appendix A Training details
Similar to [20], we slightly alter the original MPM datasets to add boundary particles, as the original data from [18] does not have them. We define the velocity at a timestep to be $\mathbf{v}_{t}=\mathbf{v}_{t}-\mathbf{v}_{t-1}$. We therefore skip the first step during training for which no velocity is available.
@@
Our implementation, training scripts, experiment configurations, and instructions for reproducing results are publicly available at [URL]. We implement NeuralMPM using PyTorch [59], and use PyTorch Geometric [60] for implementing efficient particle-to-grid functions, more specifically from the Scatter and Cluster modules. For memory efficiency, we do not store all (up to) 1,000 training trajectories in memory, and rather use a buffer of about 16 trajectories over which several epochs are performed before loading a new buffer of random trajectories.
@@
We use the official implementations and training instructions of GNS [18] and DMCF [20] to reproduce their results and conduct new experiments. More specifically, we train GNS as instructed for 5 million steps on all four datasets, using their provided configuration. For DMCF, we follow their default configurations, conducting 50k training iterations for WBC-SPH and 40k for WaterRamps, SandRamps, and Goop.
@@
We normalize the input of the model over each channel individually. We investigated computing the statistics across a buffer, resembling [61], and precomputing them on the whole training set and found no difference in performance. During inference, we use the precomputed statistics.
@@ Appendix B Supplementary results
Although we have used a U-Net architecture for the grid-to-grid processor, NeuralMPM can be used with any grid-to-grid processor and is not limited to that network. For example, in Figure 8 and Table 3 we present qualitative and quantitative ablation results, respectively, for NeuralMPM using an FNO network [51] as the grid-to-grid processor. Results show that the FNO processor is slightly worse than the U-Net processor.
Data FNO with noise FNO without noise
WaterRamps 16.8 16.3
SandRamps 5.5 3.5
Goop 4.3 3.8
Table 4 reports the exact MSE rollout values that were reported in Figure 3 for Goop.
@@
Finally, Figure 9 shows the error during rollouts for each dataset, both in terms of MSE and EMD. With both metrics, the error starts low and slowly accumulates over time. For the EMD, we use the Sinkhorn algorithm provided by [62].
Figure 9: Error propagation during rollout. Shown is the mean and 80th percentile of the MSE and EMD, computed over particles and simulations, at each timestep during the rollout. The accuracy decreases as errors accumulate.
Appendix C Dam Break 2D (LagrangeBench)
In Table 5, we report the accuracy of NeuralMPM against the two baseline checkpoints provided in LagrangeBench [63]. Note that the values reported in Table 1 for GNS were reproduced by training for 20M steps following the indications of [18], which took about 12 days using the Jax implementation of LagrangeBench. NeuralMPM, on the other hand, took 7h. In Table 6, we report the inference times of NeuralMPM and the two baselines provided in LagrangeBench. Although these are more recent implementations of graph network-based simulators, NeuralMPM still outperforms them by a large margin.
[…]
NeuralMPM (ours) 20.76 2.88
GNS [18] 114.40 224.1
SEGNN [56] 124.39 268.4
Table 5: Full-trajectory MSE (×$\times 10^{-3}$) and Sinkhorn distance (EMD) (×$\times 10^{-4}$) for NeuralMPM, GNS [18], and SEGNN [56] on the 2D Dam break dataset from LagrangeBench [63]. The two latter models are checkpoints provided by LagrangeBench. EMD is computed using the scripts provided by the authors.
One model call ($T=1$) Full rollout ($T=401$) Average FPS
NeuralMPM (ours) 7.41 193.50 2072
GNS [18] 20.46 8,170.47 49
SEGNN [56] 46.04 18,194.10 22
Table 6: Inference time (in ms) of NeuralMPM, GNS [18], and [56] on the 2D Dam break dataset from LagrangeBench [63]. Times were averaged over all test trajectories. NeuralMPM predicts 16 frames in a single model call and still outperforms the two baselines per call, which further widens the gap for the total rollout time.
Appendix D Additional predicted trajectories
In addition to the trajectories in Figures 2 and 4, we show additional trajectories emulated with NeuralMPM for all datasets in Figures 1213141517, and 16. We also release videos in the supplementary material, which we recommend watching to better see the details and limitations of NeuralMPM. This includes about 10 videos of emulated trajectories on held-out test simulations. Notably, we can observe the specific limitations of NeuralMPM on WBC-SPH, as shown in the latest trajectory depicted in Figure 16.
$t=0$ $t=75$ $t=125$ $t=200$ $t=400$ $t=997$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Initial Conditions Snapshot 1 Snapshot 2
@@ Appendix D Additional predicted trajectories
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
$t=0$ $t=119$ $t=239$ $t=358$ $t=478$ $t=598$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$t=0$ $t=79$ $t=159$ $t=238$ $t=318$ $t=398$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$t=0$ $t=32$ $t=65$ $t=97$ $t=130$ $t=398$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$t=0$ $t=119$ $t=239$ $t=358$ $t=478$ $t=598$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$t=0$ $t=100$ $t=300$ $t=500$ $t=700$ $t=3196$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: Additional WBC-SPH predicted trajectories. Snapshots of predicted trajectories against ground truth. All trajectories come from the held-out test set. To better show the differences of these longer sequences, we select the following timesteps not even in time: $t\in\{0,100,300,500,700,3199\}$. In the last trajectory, NeuralMPM struggles to follow the gravity direction and breaks down over time.
$t=0$ $t=125$ $t=400$
Ground Truth
Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption
Prediction
Refer to caption Refer to caption Refer to caption

### v2

@@ Appendix A Training details
Similar to Prantl et al. (2022), we slightly alter the original MPM datasets to add boundary particles, as the original data from Sanchez-Gonzalez et al. (2020) does not have them. We define the velocity at a timestep to be $\mathbf{v}_{t}=\mathbf{v}_{t}-\mathbf{v}_{t-1}$. We therefore skip the first step during training for which no velocity is available.
@@
Our implementation, training scripts, experiment configurations, and instructions for reproducing results are publicly available at https://neuralmpm.isach.be. We implement NeuralMPM using PyTorch (Paszke et al., 2019), and use PyTorch Geometric (Fey & Lenssen, 2019) for implementing efficient particle-to-grid functions, more specifically from the Scatter and Cluster modules. For memory efficiency, we do not store all (up to) 1,000 training trajectories in memory, and rather use a buffer of about 16 trajectories over which several epochs are performed before loading a new buffer of random trajectories.
@@
We use the official implementations and training instructions of GNS (Sanchez-Gonzalez et al., 2020) and DMCF (Prantl et al., 2022) to reproduce their results and conduct new experiments. More specifically, we train GNS as instructed for 20M steps on all four datasets, using their provided configuration. For DMCF, we follow their default configurations and train for 400K iterations for each dataset. In datasets that were not used by the original authors, VariableGravity and Dam break 2D, we performed hyperparameter search. GNS and DMCF both were trained for a total budget of 60 GPU-days per dataset, and the best performance was reported.
@@
We normalize the input of the model over each channel individually. We investigated computing the statistics across a buffer, resembling (Ioffe & Szegedy, 2015), and precomputing them on the whole training set and found no difference in performance. During inference, we use the precomputed statistics.
Code.
The code, together with additional videos, is available at the project’s website https://neuralmpm.isach.be.
@@ Appendix B Supplementary results
Evaluation.
In Table 2, we report the numerical MSE rollout values that were reported in the bar plots depicted in Figure 3 for Goop. Also, Figure 9 displays the error when rolling out a model for each dataset, both in terms of MSE and EMD. For both metrics, the error starts low and slowly accumulates over time. For the EMD, we use the Sinkhorn algorithm provided by (Cuturi et al., 2022).
@@
Figure 9: Error propagation during rollout. We show the mean and 80th percentile of the MSE and EMD, computed over particles and simulations, at each timestep during the rollout. The accuracy decreases as errors accumulate.
Grid-to-grid network.
Although we have used a U-Net architecture for the grid-to-grid processor, NeuralMPM can be used with any grid-to-grid processor and is not limited to that network. For example, in Figure 10 and Table 3 we present qualitative and quantitative ablation results, respectively, for NeuralMPM using an FNO network (Li et al., 2021) as the grid-to-grid processor. Results show that the FNO processor is slightly worse than the U-Net processor.
Data FNO with noise FNO without noise
WaterRamps 16.8 16.3
SandRamps 5.5 3.5
Goop 4.3 3.8
Offloading gravity.
Additionally, we tried adding the gravity explicitly to the velocity output in VariableGravity, but saw degraded performance compared feeding the gravity as an input to the model for large time-bundling values.
Additional inverse problem examples.
We show two additional optimization examples in Figure 11.
Initial Unoptimized Optimized
@@
Figure 11: Inverse design problem. Additional optimization examples. We exploit NeuralMPM’s differentiability to optimize the angle $\alpha$ of a ramp, anchored at the red dot, in order to get the water close to the red square region.
Appendix C Gallery of predicted trajectories
In addition to the trajectories in Figures 2 and 4, we show additional trajectories emulated with NeuralMPM for all datasets in Figures 12131415161718, and 19. We also release videos in the supplementary material, which we recommend watching to better see the details and limitations of NeuralMPM. This includes 10 videos per dataset of emulated trajectories on held-out test simulations.
Figure 17: Additional VariablyGravity predicted trajectories. Snapshots of predicted trajectories against ground truth. All trajectories come from the held-out test set.